### v1

# Rod Structures and Patching Matrices: a review

Abstract
I review the twistor theory construction of stationary and axisymmetric, Lorentzian signature solutions of the Einstein vacuum equations and the related toric Ricci-flat metrics of Riemannian signature, [36, 34, 11, 12]. The construction arises from the Ward construction [35] of anti-self-dual Yang-Mills fields as holomorphic vector bundles on twistor space, with the observation of Witten [37] that the Einstein equations for these metrics include the anti-self-dual Yang-Mills equations. The principal datum for a solution is the holomorphic patching matrix $P$ for a holomorphic vector bundle on a reduced twistor space, and $P$ is typically simpler than the corresponding metric to write down.
I give a catalogue of examples, building on earlier collections [11, 16], and consider the inverse problem: how far does the rod structure of such a metric, together with its asymptotics, determine $P$?
1 Introduction
Given the fairly recent discovery of the Chen-Teo metric [7, 8] and the recent classification of Hermitian toric ALF gravitational instantons [3], this seems a good moment to offer a review of the twistor construction of four-dimensional Ricci-flat metrics with two commuting symmetries and to collect together the known examples. A word on terminology: the method can be adapted to Lorentzian or Riemannian signature; in the Lorentzian case the symmetry is usually described as ‘stationarity and axisymmetry’ (though other two-dimensional symmetry groups may be of interest and can be handled), because the time-symmetry isn’t made periodic, while in the Riemannian case it’s always ‘toric’, i.e. with two periodic commuting symmetries.
Historically, the subject begins with Lou Witten’s paper [37] which noted that the equation governing static, axisymmetric self-dual [^2] It’s not important whether we specify self-dual or anti-self-dual here, and the former is shorter to write! $SU(2)$ Yang-Mills fields in Euclidean 4-space was the same as the main equation for stationary, axisymmetric solutions of the Einstein vacuum equations in Lorentzian 4-space. This was picked up by Ward [36] who observed that his construction [35] of solutions to self-dual Yang-Mills theory in terms of holomorphic vector bundles on (regions of) flat twistor space $\mathbb{CP}^{3}$, could therefore be applied to generate these Einstein vacua, giving solutions in terms of holomorphic vector bundles over a twistor space reduced by the symmetry.
The next step was taken by Mason and Woodhouse [34]. They gave a detailed account of the reduced twistor space, which is essentially the twistor space $\mathbb{CP}^{3}$ or a large region in it, quotiented by two holomorphic vector fields derived from the two space-time symmetries. The action is not free and proper and the reduced twistor space $R$ turns out to be a one-dimensional non-Hausdorff complex manifold, essentially two copies of the Riemann sphere $\mathbb{CP}^{1}$ with identifications. It’s still the case that holomorphic vector bundles $E$ on $R$ can be defined; this is done most conveniently with respect to a fixed standard four-set cover of $R$, two sets for each $\mathbb{CP}^{1}$, and then just one of the transition matrices has the information of the space-time in it – call this the patching matrix $P$ – and the information of the metric can be extracted by splitting $P$ according to Ward’s prescription [36]. Different reality conditions lead to real Riemannian or real Lorentzian solutions.
Fletcher and Woodhouse [11, 12] gave more examples and considered the inverse problem: without looking at the metric, what data from the space-time determine $P$ and therefore the metric? In this way, the Lorentzian Kerr metric is determined by its rod structure (defined below) and its asymptotic conserved quantities. Clearly this gives another route to black hole uniqueness, which was their aim, and emphasises the status of $P$ as the fundamental datum for the metric. Finally, I should mention the extension of the method to five-dimensional black holes with three commuting symmetries due to Metzner and coworkers [25, 32] and the catalogue of examples compiled by Gray [16].
The contents of this article are as follows: in Section 2, I briefly review the theory of stationary, axisymmetric space-times and give a sketch of the twistor theory described above; the notion of rod structure of a solution is discussed in Section 3; the main section is 4 where a catalogue of examples of $P$ is presented, some original and some collected from earlier work; finally there is a discussion of the inverse problem, the problem of obtaining $P$ directly from the rod structure and asymptotic quantities, in Section 5. Some extra material is gathered in two appendices.
Acknowledgements
I gratefully acknowledge assistance from Maciej Dunajski, Lionel Mason and Nick Woodhouse in preparing this review and catalogue, and to Lars Andersson for the invitation to speak about this construction at a meeting at the Mittag-Leffler Institute in July 2023, the preparation for which prompted this review.
2 Stationary, axisymmetric vacua and their twistor theory
2.1 Stationary axisymmetric vacua
We start by reviewing the Einstein vacuum equations for stationary, axisymmetric, Lorentzian solutions, following [29], [34] and [36]. Suppose the two Killing vectors are $K=\partial/\partial t$ for time-translation and $L=\partial/\partial\phi$ for axisymmetry and that these commute. Also that the isometry group is orthogonally transitive, that is the 2-surfaces orthogonal to the suraces of transitivity are integrable. For vacuum solutions the two scalars $*(K\wedge L\wedge dK)$ and $*(K\wedge L\wedge dL)$ are necessarily constant and they will vanish if either $K$ or $L$ has a fixed point (which $L$ usually does have); then their common vanishing is the condition for orthogonal transitivity, which in turn ensures that coordinates can be chosen so that the metric is block $2\times 2$.
Now we can write the metric in the form
$g=h_{ij}dx^{i}dx^{j}+\left(\begin{array}[]{cc}dt&d\phi\\ \end{array}\right)J\left(\begin{array}[]{rr}dt\\ d\phi\\ \end{array}\right)$ (1)
with $i,j=1,2$ and $J$, sometimes called the Gram or Yang matrix, is
$J=\left(\begin{array}[]{cc}g_{tt}&g_{t\phi}\\ g_{t\phi}&g_{\phi\phi}\\ \end{array}\right)=\left(\begin{array}[]{cc}U&V\\ V&W\\ \end{array}\right)\mbox{ say.}$.
With Lorentzian signature (+)(+---)( + - - - ), $h_{ij}$ is negative definite and $J$ has negative determinant. We can include Riemannian metrics in the ansatz (1) by taking $h_{ij}$ and $J$ to be positive definite. In this case usually both Killing vectors have closed trajectories and the metric is said to be toric (because of its torus symmetry).
[…]
or in the Riemannian case by
[…]
In either case, we choose $z$ to be the harmonic conjugate of $r$ (with respect to the 2-metric $h_{ij}$) and then the metric $h_{ij}$ becomes
$h_{ij}dx^{i}dx^{j}=\pm\Omega^{2}(dr^{2}+dz^{2}),$ (4)
with plus for Riemannian and minus for Lorentzian. Note that $r$ is uniquely specified up to sign and $z$ up to sign and an additive constant. The locus of vanishing $r$, which we consider in detail in Section 3, is clearly of crucial interest and can be thought of for now as the axis of symmetry. The coordinates $(r,z)$ thus defined are often referred to as Weyl-Papapetrou coordinates.
It’s quite common in the literature to write out (1) at greater length in the Lorentzian case, in what is also called the Weyl-Papapetrou form, as
$g=-f^{-1}(e^{2k}(dr^{2}+dz^{2})+r^{2}d\phi^{2})+f(dt-\omega d\phi)^{2},$ (5)
when $\Omega^{2}=f^{-1}e^{2k}$ and then [12], [34]
[…]
and $\mbox{det}J=-r^{2}$ so that the $r,z$ in (5) are the Weyl-Papapetrou coordinates. The vacuum field equations for the metric (5) can be found (with minor changes of convention) in [29] (which can be taken as the reference for this part) as
$\Delta(\log f)$ $=$ $-\frac{f^{2}}{r^{2}}(\omega_{r}^{2}+\omega_{z}^{2})$ (7)
$\left(\frac{f^{2}\omega_{r}}{r}\right)_{r}+\left(\frac{f^{2}\omega_{z}}{r}\right)_{z}$ $=$ 00 0 (8)
$k_{r}$ $=$ $\frac{r}{4f^{2}}(f_{r}^{2}-f_{z}^{2})-\frac{f^{2}}{4r}(\omega_{r}^{2}-\omega_{z}^{2})$ (9)
$k_{z}$ $=$ $\frac{r}{2f^{2}}f_{r}f_{z}-\frac{f^{2}}{2r}\omega_{r}\omega_{z}$ (10)
where
$\Delta F(r,z)=F_{rr}+\frac{1}{r}F_{r}+F_{zz},$
which is the 3-dimensional flat-space Laplacian in cylindrical polar coordinates for an axisymmetric function $F$. The integrability condition for (9, 10) is satisfied by virtue of (7, 8)
Remarkably, (7) and (8) can be written as the matrix equation
$r^{-1}\partial_{r}(rJ^{-1}\partial_{r}J)+\partial_{z}(J^{-1}\partial_{z}J)=0,$ (11)
[37], usually called Yang’s equation, and then the equations (9, 10) can be written as a single (complex) equation for $\Omega$:
$2i\partial_{w}\log(r\Omega^{2})=r\mbox{Tr}(-J^{-1}J_{w}J^{-1}J_{w}),$ (12)
where $w=z+ir$ and Tr means trace. As is to be expected, the integrability condition for (12) is satisfied by virtue of (11). It’s worth noting that there is the freedom to add a constant of integration to $\log(r\Omega^{2})$, which translates as a free multiplicative constant on $\Omega$. This can be thought of as the origin of the constant $k$ in (71).
We’ll be concentrating on (11), but $\Omega$ and therefore (12) are necessary for a discussion of conical singularities on the axis, which we’ll be paying less attention to.
The next step towards solving the field equations has traditionally been, since [10], to introduce a potential for $\omega$: (8) is the integrability condition for the existence of a function $\psi$ satisfying
$\psi_{r}=\frac{f^{2}}{r}\omega_{z},\;\;\psi_{z}=-\frac{f^{2}}{r}\omega_{r},$
and fixed only up to an additive constant. These equations for $\psi$ turn out to be equivalent to
$d\psi=*(K\wedge dK),$
with $K=\partial_{t}$ as before. Thus $\psi$ is often referred to as the twist potential for the Killing vector $K$, since $*(K\wedge dK)$ can be called the twist of the Killing vector, and this is co-closed by virtue of the vacuum equations. Evidently one could define a different twist potential by starting with a different Killing vector, and we’ll take this up later (see (55)). However, starting with $K$ we may consider the matrix [12], [34], [37]
[…]
then $J^{\prime}$ also satisfies (11) but now, note, with unit determinant.The replacement of $J$ by $J^{\prime}$ can be regarded as a Bäcklund transformation of the Yang equation (11) (see [34] or [25]). Note that one may recover $J$ from a knowledge of $J^{\prime}$, and that there will be a variety of different $J^{\prime}$ for the same $J$, obtained from different choices of Killing vector.
It’s worth noting that there is a version of (12) involving only $J^{\prime}$, namely
$4ik_{w}=Tr(-J^{\prime-1}J^{\prime}_{w}J^{\prime-1}J^{\prime}_{w}),$ (14)
with $k$ as in (5).
Following [12], we shall call the matrix $J^{\prime}$ the Ernst potential for $J$ determined by $K$ (the more familiar scalar Ernst potential is the function $f+i\psi$, see e.g. [29]).
So far we have considered Lorentzian metrics. The simplest way to shift to Riemannian metrics is to replace $J$ in (6) by [^3] This is done in [16] and seems to be due originally to Mason.
[…]
$g=f^{-1}(e^{2k}(dr^{2}+dz^{2})+r^{2}d\phi^{2})+f(dt-\omega d\phi)^{2}.$ (16)
Remarkably, (11) and (12) are both unchanged, but, with $\psi$ defined as before, $J^{\prime}$ becomes
[…]
now with determinant -1.
2.2 Twistor construction of stationary axisymmetric vacuum solutions
It was Witten’s observation [37] that Yang’s equation (11) also determines anti-self-dual, static [^4] In the literature of relativity ‘stationary’ means possessing a time-like Killing vector, while ‘static’ means possessing a hypersurface-orthogonal time-like Killing vector., axisymmetric gauge fields on $\mathbb{CM}$, and that these are solved by the Ward construction [35], so we need a quick sketch of that. An anti- self-dual Yang-Mills field on the complexification of (a region of) Minkowski space $\mathbb{CM}$ defines a connection $\mathcal{A}$ which is flat on $\alpha$-surfaces i.e. on self-dual 2-planes. It therefore defines a vector space of constant sections on each $\alpha$-plane, and hence a holomorphic vector bundle $\mathcal{E}$ on the corresponding region in the space of $\alpha$-planes, which is projective twistor space $\mathbb{PT}=\mathbb{CP}^{3}$. Conversely $\mathcal{E}$ determines $\mathcal{A}$: restrict $\mathcal{E}$ to a projective line $\mathbb{CP}^{1}$ in $\mathbb{PT}$; it’s part of standard twistor theory that these lines represent points in $\mathbb{CM}$ so by its construction $\mathcal{E}$ is trivial on these lines, and so this restriction splits; again it’s part of standard twistor theory that the splitting determines $\mathcal{A}$ at the corresponding point in $\mathbb{CM}$. Reality conditions on $\mathcal{A}$ can be straightforwardly imposed.
If $\mathcal{A}$ is constant along a Killing vector $K$ of $\mathbb{CM}$, then $\mathcal{E}$ is constant along a corresponding holomorphic vector field on $\mathbb{CP}^{3}$ and one can contemplate taking quotients. This can similarly be done with two commuting symmetries, but now as Mason and Wodehouse made clear, the action is not free and proper, so that the quotient, though a one-dimensional complex manifold, is not Hausdorff. However one may boldly go on and consider vector bundles on this non-Hausdorff Riemann surface. The pay-off is that, associated with a rod, that is to say a segment of the symmetry axis $r=0$, on which the Killing vector $L$ vanishes and the Killing vector $K$ does not, the non-trivial patching matrix $P(z)$ of the bundle is obtained from the Ernst potential $J^{\prime}(r,z)$ corresponding to $K$ simply by setting $r=0$:
$P(z)=J^{\prime}(0,z).$ (18)
It follows that there is a good deal of freedom in $P$, for example from changing the choice of $K$ or by adding a constant to $\psi$, which corresponds to conjugation of $P$ with a constant matrix.
A good deal more could be said here, but the original publications are clear and readily available.
3 Rod structures
The reference for this section is [18].
The locus $r=0$ in the $(r,z)$-half-plane from the previous section can loosely be regarded as the axis of symmetry, and transformed into a straight line (by the Riemann Mapping Theorem if necessary). It denotes the set of points where the rank of $J$ is strictly less than two, and contains a certain number of nodes (also called nuts in the literature, which risks confusion with nuts in the classification of [14], or turning points - see [8]) where the rank is actually zero, connected by rods on which the rank is one. The location of these nodes (in the $z$-coordinate) together with a label on each rod indicating which Killing vector lies in the kernel of $J$ there is what is called the rod structure of the solution [18]. Together with asymptotic quantities, usually the mass and angular momentum, in the ALF case the rod structure characterises a solution (proved for five-dimensional AF Lorentzian in [20]; the case of four-dimensional AF Lorentzian is ‘black hole uniqueness’ and in this context we can refer to [23]). Call this set of objects the data, then there’s no known prescription for generating all and only those data which are compatible with a solution. We’ll give some examples.
3.1 $\mathbb{E}^{4}$
By this we mean flat Riemannian (i.e. Euclidean) four-space with the symmetry group chosen as two commuting rotations, in orthogonal planes (other choices are of course possible). Write the metric in terms of two sets of plane polars:
$g=dR_{1}^{2}+dR_{2}^{2}+\left(\begin{array}[]{cc}d\phi_{1}&d\phi_{2}\\ \end{array}\right)\left(\begin{array}[]{cc}R_{1}^{2}&0\\ 0&R_{2}^{2}\\ \end{array}\right)\left(\begin{array}[]{rr}d\phi_{1}\\ d\phi_{2}\\ \end{array}\right),$ (19)
then $r^{2}=\mbox{det}J=R_{1}^{2}R_{2}^{2}$ and so w.l.o.g. $r=R_{1}R_{2}$, when w.l.o.g. $z=\frac{1}{2}(R_{1}^{2}-R_{2}^{2})$. The axis of symmetry is the line $r=0$ and it has a node at the origin. The upper rod is at $R_{2}=0$, where $\partial_{\phi_{2}}$ vanishes and $z=\frac{1}{2}R_{1}^{2}$ is positive; the lower rod is at $R_{1}=0$ where $\partial_{\phi_{1}}$ vanishes and $z=-\frac{1}{2}R_{2}^{2}$ is negative. It’s rather clear that there could be no nonsingular solution with this rod structure and positive mass, which serves to show that some data are incompatible with the existence of a solution.
This example also typifies more complicated but asymptotically Euclidean (AE) or asymptotically locally Euclidean (ALE) solutions. Note the outermost rods are defined by a pair of linearly independent Killing vectors, which therefore automatically provide a basis of Killing vectors.
3.2 The Schwarzschild solution
For the Lorentzian Schwarzschild metric in the usual coordinates, $J$ is given by
$J=\left(\begin{array}[]{cc}1-2m/R&0\\ 0&-R^{2}\sin^{2}\theta\\ \end{array}\right)$
and we’ll take $m>0$. One speedily obtains
$r=\sqrt{R(R-2m)}\sin\theta,\;z=(R-m)\cos\theta,$
from which the rod structure can be read off by setting $r=0$: there are 3 rods, respectively at (i) $\theta=0$, (ii)$R=2m$ and (iii) $\theta=\pi$, with nodes at $z=\pm m$. The central rod is the horizon, but in Riemannian Schwarzschild is the prime example of a bolt in the classification of [14], being a 2-sphere fixed by the isometry group.
For the top rod we have $\theta=0$ and so $z=R-m$, and since the Killing vector $\partial/\partial t$ is hypersurface-orthogonal we may take $\psi=0$. Thus the patching matrix $P_{+}$ from (13) (and given already in [11], [12] and [36]) is
$P_{+}=\left(\begin{array}[]{cc}\frac{z+m}{z-m}&0\\ 0&\frac{z-m}{z+m}\\ \end{array}\right).$
We can use the algorithm from the next section for going past a node, to obtain the patching matrices for all rods as
$P_{0}=\left(\begin{array}[]{cc}4(z^{2}-m^{2})&0\\ 0&(4(z^{2}-m^{2}))^{-1}\\ \end{array}\right)$
$P_{-}=\left(\begin{array}[]{cc}\frac{z-m}{z+m}&0\\ 0&\frac{z+m}{z-m}\\ \end{array}\right).$
For Riemannian Schwarzschild flip the sign on the lower right term in each $P_{i}$.
3.3 Information from asymptotics
The Schwarzschild solution, Lorentzian or Riemannian, is asymptotically flat and so has the asymptotics identified in [18] for stationary, axisymmetric Ricci-flat metrics (though in the Lorentzian case these must have been written down earlier). In terms of a spherical radial coordinate $R$ these have the expansion (see also [2] or [23])
$U=1-\frac{2m}{R}+O(R^{-2}),\;\;V=\frac{2L}{R}\sin^{2}\theta(1+O(R^{-1})),\;\;W=R^{2}\sin^{2}\theta(1+O(R^{-1})),$
with $m,L$ the total mass and angular momentum (and this is now the Riemannian case). The Weyl-Papapetrou coordinates are
$r=(R(R-2m))^{1/2}\sin\theta(1+O(R^{-1})),\;\;z=(R-m)\cos\theta(1+O(R^{-1})),$
and the twist potential for the Killing vector $K=\partial_{t}$ is
$\psi=\frac{2L}{R^{2}}\cos\theta(1+O(R^{-1})).$
The outer rods are immediately identifiable as $\theta=0,\pi$. For the top rod at $\theta=0$ we obtain
$P_{+}=\left(\begin{array}[]{cc}1+\frac{2m}{z}+O(z^{-2})&-\frac{2L}{z^{2}}+O(z^{-3})\\ -\frac{2L}{z^{2}}+O(z^{-3})&-1+\frac{2m}{z}+O(z^{-2})\\ \end{array}\right).$ (20)
This expansion will be useful below. In particular, it shows that for Riemannian ALF we may suppose wlog that $P_{+}\rightarrow\left(\begin{array}[]{cc}1&0\\ 0&-1\\ \end{array}\right)$ as $z\rightarrow\infty$. This was shown in [11] for the Lorentzian case when instead $P_{+}\rightarrow I$.
Having obtained the asymptotics for AF/ALF, it’s worth taking a moment to do the corresponding thing for AE/ALE. Following a suggestion of James Lucietti, we can obtain the asymptotics of 4-dimensional AE/ALE from those implicit in [18] for those particular 5-dimensional AF/ALF vacua which are just 4-dimensional vacua times a line. We consider the perturbation of (19), changing the coordinates to eliminate subscripts:
$g=\Theta^{2}(du^{2}+dv^{2})+\left(\begin{array}[]{cc}d\phi&d\chi\\ \end{array}\right)\left(\begin{array}[]{cc}u^{2}(1+\Phi)&\Xi\\ \Xi&v^{2}(1-\Phi)\\ \end{array}\right)\left(\begin{array}[]{rr}d\phi\\ d\chi\\ \end{array}\right),$ (21)
where $\Phi,\Xi$ are to be regarded as small, and $\Theta$ is close to one (we’ve used a gauge transformation to set the trace of the metric perturbation to zero). For the Weyl-Papapetrou coordinates we obtain
$r=uv(1+h.o.),\;\;z=\frac{1}{2}(u^{2}-v^{2}+h.o.).$
Linearising the Yang equation (11) with this $J$ leads to the Laplace equation for $\Phi$
$(uv\Phi_{u})_{u}+(uv\Phi_{v})_{v}=0,$ (22)
and an equation for $\Xi$:
$(\frac{1}{uv}\Xi_{u})_{u}+(\frac{1}{uv}\Xi_{v})_{v}=0.$ (23)
This last is simplified by the substitution
$S_{v}=\frac{1}{uv}\Xi_{u},\;\;S_{u}=-\frac{1}{uv}\Xi_{v},$
when it becomes the Laplace equation for $S$:
$(uvS_{u})_{u}+(uvS_{v})_{v}=0.$
We choose a monopole solution for $\Phi$ and a symmetric dipole-like solution for $S$:
$\Phi=\frac{2M}{(u^{2}+v^{2})},\;\;S=\frac{L(u^{2}-v^{2})}{(u^{2}+v^{2})^{3}},$ (24)
when
$\Xi=\frac{2Lu^{2}v^{2}}{(u^{2}+v^{2})^{3}}.$ (25)
With these substituted into (21) we obtain the asymptotics given in [18], with $M,L$ here related to $\eta,\zeta$ there [^5] It may be misleading to call these quantities $M$ and $L$ – a more neutral usage would be $\eta,\zeta$ but, up to constant factors, they do occupy the places filled by $M,L$ in the AF/ALF cases.
For the corresponding $P$, we take the top rod to be $v=0$ so that the Killing vector $\partial_{\chi}$ vanishes there, and compute the twist potential of the Killing vector $\partial_{\phi}$ to be
$\psi=-\frac{2Lu^{4}}{(u^{2}+v^{2})^{3}}.$
On the axis where $v=0$, so $z=u^{2}/2$, we obtain
$P_{+}=\left(\begin{array}[]{cc}\frac{1}{2z}\left(1-\frac{M}{z}+h.o.\right)&\frac{L}{2z^{2}}+h.o.\\ \frac{L}{2z^{2}}+h.o.&-2z(1+\frac{M}{z}+h.o.)\\ \end{array}\right),$ (26)
which we’ll use below.
3.4 Patching matrices from rod structure and $J$
This is the starting point for the inverse problem. I’ll quote a result from [11] for an operation conveniently called ‘passing by a node’. Suppose then we have a rod structure with $k$ nodes located at $z=a_{i},i=1,\ldots,k$ and a basis $(K,L)$ of Killing vectors. We know $J$ and on a rod on which $K$ does not vanish we can calculate $J^{\prime}$. This will be singular at the next node, but it will at most have simple poles there (otherwise there would be a degenerate horizon in the Lorentzian case or a singularity in the Riemannian case) and the procedure for passing this node is as follows: suppose the node is at $z=0$ then there is freedom $P\rightarrow CPC^{T}$ with constant $C$ of unit determinant to set
$P=P^{+}=\left(\begin{array}[]{cc}(2z)^{-1}a(z)&b(z)\\ b(z)&(2z)c(z)\\ \end{array}\right)\mbox{ with holomorphic }a,b,c$ (27)
above the node, when below the node we set
$P=P^{-}=\left(\begin{array}[]{cc}(2z)a(z)&b(z)\\ b(z)&(2z)^{-1}c(z)\\ \end{array}\right).$
From this prescription for passing by a node, we see that poles can only be introduced or removed at a node, so in general the entries in $P(z)$ can only be rational functions with simple poles at the nodes (poles off the axis give rise to singularities in the space-time metric – see Appendix 2). If the nodes are located at $z=a_{i}$ for $i=1,\ldots,k$ then a priori the denominators of all entries in $P$ are $D(z)=\Pi_{i}^{k}(z-a_{i})$ and numerators are polynomials, furthermore with their degrees fixed by the asymptotics. Also in AF/ALF or AE/ALE cases some coefficients in a series expansion are fixed by the asymptotic quantities. The determinant condition on $P(z)$ then fixes more, and in simple cases all, of the remaining coefficients. We’ll return to this question in Section 5.
4 Examples
Now we present a menagerie of examples, paying particular attention to the Hermitian ones because of [3].
4.1 $\mathbb{E}^{4}$
As we saw above, on the upper rod $z=\frac{1}{2}R_{1}^{2}$ and the metric component $g_{\phi_{1}\phi_{1}}$ is $R_{1}^{2}=2z$ so from (17) the patching matrix for the upper rod is
$P_{+}=\left(\begin{array}[]{cc}(2z)^{-1}&0\\ 0&-(2z)\\ \end{array}\right).$ (28)
Following the same algorithm for $P_{-}$ gives
$P_{-}=\left(\begin{array}[]{cc}-(2z)^{-1}&0\\ 0&(2z)\\ \end{array}\right),$
while taking $P_{+}$ past the node at $z=0$ would give
$\widetilde{P}_{-}=\left(\begin{array}[]{cc}2z&0\\ 0&-(2z)^{-1}\\ \end{array}\right),$
which we regard as equivalent, since
$P_{-}=\left(\begin{array}[]{cc}0&1\\ 1&0\\ \end{array}\right)\widetilde{P}_{-}\left(\begin{array}[]{cc}0&1\\ 1&0\\ \end{array}\right).$
There are obviously no conicality issues in this case, and the explicit $P_{+}$ raises the expectation that generally for AE/ALE, $P_{+}$ does not have a limit as $z\rightarrow\infty$ but rather has terms with a simple zero and terms with a simple pole. We’ll see an example like this with the C-metric in Section 4.7.
4.2 Kerr
The Lorentzian Kerr metric taken from [19] but with signature switched to (+)(+---)( + - - - ) is
$g_{L}=\left(1-\frac{2mR}{\Sigma^{2}}\right)dt^{2}+\frac{4maR\sin^{2}\theta}{\Sigma^{2}}d\phi dt-\left((R^{2}+a^{2})\sin^{2}\theta+\frac{2mRa^{2}}{\Sigma^{2}}\sin^{4}\theta\right)d\phi^{2}-\Sigma^{2}\left(\frac{dR^{2}}{\Delta}+d\theta^{2}\right),$ (29)
in terms of two constants $m,a$ and with $\Sigma^{2}=R^{2}+a^{2}\cos^{2}\theta$ and $\Delta=R^{2}-2mR+a^{2}$, and is asymptotically-flat (AF). In the Lorentzian case, we’ll require $m^{2}>a^{2}$ to avoid having a degenerate horizon, but we shall mainly be concerned with the Riemannian form of the metric, obtained by the substitution $(t,a)\rightarrow(it,ia)$ and a switch of overall sign (the Lorentzian case, up to the derivation of the patching matrices $P_{i}$, is given in [11], [12]). This is
$g_{R}=\left(1-\frac{2mR}{\Sigma^{2}}\right)dt^{2}+\frac{4maR\sin^{2}\theta}{\Sigma^{2}}d\phi dt+\left((R^{2}-a^{2})\sin^{2}\theta-\frac{2mRa^{2}}{\Sigma^{2}}\sin^{4}\theta\right)d\phi^{2}+\Sigma^{2}\left(\frac{dR^{2}}{\Delta}+d\theta^{2}\right),$ (30)
where now $\Sigma^{2}=R^{2}-a^{2}\cos^{2}\theta$ and $\Delta=R^{2}-2mR-a^{2}$, and the metric is AF.
In the terminology of Section 2, we have
$J=\left(\begin{array}[]{cc}1-\frac{2mR}{\Sigma^{2}}&\frac{2maR\sin^{2}\theta}{\Sigma^{2}}\\ \frac{2maR\sin^{2}\theta}{\Sigma^{2}}&(R^{2}-a^{2})\sin^{2}\theta-\frac{2ma^{2}R}{\Sigma^{2}}\sin^{4}\theta\\ \end{array}\right),$ (31)
and then
$r=(\mbox{det}J)^{1/2}=(R^{2}-2mR-a^{2})^{1/2}\sin\theta,$ (32)
from which one rapidly finds
$z=(R-m)\cos\theta,$ (33)
having made a convenient choice of additive constant in $z$.
From (32) one rapidly identifies the rods: there are three, respectively (i) at $\theta=0$, (ii) at $R=R_{+}$ and (iii) at $\theta=\pi$, defining two nodes at $R=R_{+}$ and $\theta=0$ or $\pi$, so also at $z=\pm(m^{2}+a^{2})^{1/2}$. At the top and bottom rods, the Killing vector $L=\partial_{\phi}$ vanishes; at the middle rod the Killing vector $\partial_{t}-\frac{a}{2mR_{+}}\partial_{\phi}$ vanishes (this is in the Riemannian case when the middle rod is a bolt in the terminology of [14]; in the Lorentzian case this is the location of the outermost Killing horizon where the Killing vector $\partial_{t}+\frac{a}{2mR_{+}}\partial_{\phi}$ is null).
Adapting to the top rod, we’ll calculate the twist potential for the Killing vector $K=\partial_{t}$. From the definition
$d\psi=*(K\wedge dK),$
we find
$\psi=\frac{2ma}{R^{2}-a^{2}\cos^{2}\theta}.$ (34)
For the top rod, by evaluating on the axis, we find $R=z+m$ and [^6] This expression for $J^{\prime}$, which may be compared with the expansion in (20), differs from the expression in [11, 12, 34] as we are considering the Riemannian i.e. Wick-rotated, Kerr.
$P_{+}=J^{\prime}(0,z)=f^{-1}\left(\begin{array}[]{cc}1&-\psi\\ -\psi&\psi^{2}-f^{2}\\ \end{array}\right)=\frac{1}{(z^{2}-\sigma^{2})}\left(\begin{array}[]{cc}(z+m)^{2}-a^{2}&-2ma\\ -2ma&-(z-m)^{2}+a^{2}\\ \end{array}\right)$ (35)
where $\sigma=+(m^{2}+a^{2})^{1/2}$ and the top node is at $z=\sigma$.
In the corresponding way, we find on the lower rod
$P_{-}=J^{\prime}(0,z)=\frac{1}{(z^{2}-\sigma^{2})}\left(\begin{array}[]{cc}(z-m)^{2}-a^{2}&-2ma\\ -2ma&-(z+m)^{2}+a^{2}\\ \end{array}\right).$ (36)
Note that we could have obtained $P_{-}$ from $P_{+}$ by remarking that the Kerr metric has a discrete isometry $\theta\rightarrow-\theta$ or $z\rightarrow-z$.
For the middle rod, the matrix $P_{+}$ in (35) is not in the form of (27) for passing by the top node but if we take
$C=\left(\begin{array}[]{cc}1&0\\ \gamma&1\\ \end{array}\right)$
with $\gamma=a/(m+\sigma)$ then
$P_{+}\rightarrow CP_{+}C^{T}=\left(\begin{array}[]{cc}\frac{(z+m)^{2}-a^{2}}{(z+\sigma)(z-\sigma)}&\frac{a(z+2m+\sigma)}{(m+\sigma)(z+\sigma)}\\ \frac{a(z+2m+\sigma)}{(m+\sigma)(z+\sigma)}&\frac{-2m(z-\sigma)}{(m+\sigma)(z+\sigma)}\\ \end{array}\right),$
which now is in the right form to apply (27) and obtain $P_{0}$.
4.2.1 Hermiticity of Kerr
Since both the self-dual and the anti-self-dual Weyl tensors are type D, the Riemannian Kerr metric has two distinct Hermitian structures, one associated with each orientation. The details can be found in [31].
4.2.2 Conicality issues
There are known to be no conicality problems in the Lorentzian case: the central rod is the black-hole horizon and extension through this is well-known. For the Riemannian case see the discussion in [14].
4.3 Taub-NUT
This metric is discussed at some length in [19] and the elucidation of its structure was influential in the development of mathematical general relativity. It was first found as what would now be called a vacuum spatially homogeneous cosmology of Bianchi-type IX, [30]. The metric form becomes singular where the surfaces of homogeneity become null but there exist extensions across these surfaces, where the character of the solution changes and it was seen to be identical with a solution independently found in [26]. The standard name for the solution is obtained from these two references ([26] gives NUT=Newman-Unti-Tamburino), but the term nut in the classification of [14] was motivated by this name. We shall take the Lorentzian form of the metric from [19], with the signature switched to (+)(+---)( + - - - ) and a slight notational change, as
$g_{L}=\frac{dt^{2}}{A(t)}-A(t)(d\chi+2\ell\cos\theta d\phi)^{2}-(t^{2}+\ell^{2})(d\theta^{2}+\sin^{2}\theta d\phi^{2}),$ (37)
in terms of two constants $m,\ell$ and where $A(t)=-(t^{2}-2mt-\ell^{2})(t^{2}+\ell^{2})^{-1}$. Factorise the numerator of $A(t)$ as $(t_{+}-t)(t-t_{-})$ with $t_{\pm}=m\pm(m^{2}+\ell^{2})^{1/2}$ then for $t_{-}<t<t_{+}$ this is a Lorentzian Bianchi IX metric (the Taub solution) but the surfaces of homogeneity become null at $t=t_{\pm}$. The extensions through these surfaces, to the NUT part of the solution where $t$ becomes a radial coordinate, are discussed in [19], but we will be concerned with the Riemannian form of the metric which is obtained by the replacements $(t,\chi,\ell)\rightarrow(R,i\chi,iN)$ and a switch of overall sign (I’m using $N$ rather than $n$ to avoid confusion with a later reference to integer $n$). This is
$g_{R}=\frac{dR^{2}}{U(R)}+U(R)(d\chi+2N\cos\theta d\phi)^{2}+(R^{2}-N^{2})(d\theta^{2}+\sin^{2}\theta d\phi^{2}),$ (38)
now with $U=(R^{2}-2mR+N^{2})(R^{2}-N^{2})^{-1}=-A$. In the terminology of Section 2, we have
$J=\left(\begin{array}[]{cc}U&2NU\cos\theta\\ 2NU\cos\theta&4N^{2}U\cos^{2}\theta+(R^{2}-N^{2})\sin^{2}\theta\\ \end{array}\right),$ (39)
and then
$r=(\mbox{det}J)^{1/2}=(R^{2}-2mR+N^{2})^{1/2}\sin\theta,$ (40)
from which one rapidly finds
$z=(R-m)\cos\theta,$ (41)
with a convenient choice of additive constant in $z$. Now there are evidently three rods: call them respectively (i) $\theta=0$, (ii) $R=R_{+}=m+(m^{2}-N^{2})^{1/2}$ and (iii) $\theta=\pi,$ and two nodes at respectively $(R,\theta)=(R_{+},0)$ so $z=(m^{2}-N^{2})^{1/2}$ and $(R,\theta)=(R_{+},\pi)$ so $z=-(m^{2}-N^{2})^{1/2}$. We’ll assume $m^{2}>N^{2}>0$ to keep the nodes real and separate, and it’s convenient to write $\sigma=(m^{2}-N^{2})^{1/2}$, so the nodes are at $z=\pm\sigma$. One can think of the parameter $m$ as a mass and then $N$ is usually called the nut parameter (and is sometimes called the nuttiness).
By inspection of (39) we can see that the kernel of $J$ is $\partial_{\phi}-2N\partial_{\chi}$ on (i), $\partial_{\chi}$ on (ii) and $\partial_{\phi}+2N\partial_{\chi}$ on (iii). This should be compared with Kerr in the previous section: both have three rods and two nodes but the Killing vector labelling is quite different, so the rod structures are different, and in particular this metric won’t be AF or ALF.
We’ll calculate the twist potential for the Killing vector $K=\partial_{\chi}$. We find
$d\psi=*(K\wedge dK)=2n\frac{(R^{2}-2mR+N^{2})}{(R^{2}-N^{2})^{2}}dR,$
whence
$\psi=\frac{n(R-R_{+})^{2}}{R_{+}(R^{2}-N^{2})},$ (42)
where we’ve chosen the additive constant in $\psi$ so that $\psi(R_{+})=0$. For the top rod we set $\theta=0$ so $z=R-m$ and substituting into (17) we obtain
$P_{+}=J^{\prime}(0,z)=\left(\begin{array}[]{cc}\frac{(z+m)^{2}-N^{2}}{(z-\sigma)(z+\sigma)}&-\frac{N(z-\sigma)}{R_{+}(z+\sigma)}\\ -\frac{N(z-\sigma)}{R_{+}(z+\sigma)}&-\frac{2\sigma(z-\sigma)}{R_{+}(z+\sigma)}\\ \end{array}\right).$ (43)
The patching matrices for the other rods are straightforward to obtain by the method of subsection 3.4.
Note that $P_{+}$ in (43) tends to a constant as $z\rightarrow\infty$ but doesn’t tend to the simple form $\left(\begin{array}[]{cc}1&0\\ 0&-1\\ \end{array}\right)$ that one expects for AF or ALF. We can impose this limit by replacing $P_{+}$ by $\tilde{P}_{+}=C^{T}P_{+}C$ with
$C=\left(\begin{array}[]{cc}1&\frac{N}{R_{+}}\\ 0&1\\ \end{array}\right).$
This transformation is equivalent to adding a constant to $\psi$. It evidently preserves the symmetry and the determinant of $P_{+}$ and gives an equivalent twistor construction. We find [^7] Previously as given in [34].
$\tilde{P}_{+}=\frac{1}{(z^{2}-\sigma^{2})}\left(\begin{array}[]{cc}(z+m)^{2}-N^{2}&2Nz\\ 2Nz&-(z-m)^{2}+N^{2})\\ \end{array}\right).$ (44)
This should be compared with (36): the nut-parameter term, in the off-diagonal slot, here appears with a higher power of $z$ than the angular momentum term in (36), though it still goes to zero at large $z$.
4.3.1 Self-dual Taub-NUT and Taub-bolt
Self-dual Taub-NUT is the particular case $m=N$, so $R_{+}=m$ and $\sigma=0$: the Weyl tensor becomes self-dual and the two nodes apparently merge, but we don’t get a second-order pole in $P_{+}$, rather from (43) we find
$P_{+}=\left(\begin{array}[]{cc}1+\frac{2m}{z}&-1\\ -1&0\\ \end{array}\right),$ (45)
so that there is a single node, and it is located at the origin (apparently resembling $\mathbb{E}^{4}$, but the asymptotics are quite different, as is $P$). We’ll see below that this form of $P_{+}$ is characteristic of the Gibbons-Hawking metrics, which include self-dual Taub-NUT. There are now only two rods and on the lower one we can directly calculate the patching matrix as
$P_{-}=\left(\begin{array}[]{cc}1-\frac{2m}{z}&-1\\ -1&0\\ \end{array}\right).$ (46)
Clearly we need a different rule from the one in section 3.4 for going past a node with the Gibbons-Hawking metrics. We’ll return to this in section 4.5.
The so-called Taub-bolt metric was spotted as another particular case of the general Taub-NUT in [27]. It is neither self-dual nor anti-self-dual, but is Hermitian as we’ll see below. To obtain it, one sets $m=5N/4$, so $R_{+}=2N$ and it’s convenient to work with $\sigma$ which is now $43N/$ 4. We find
$P_{+}=\left(\begin{array}[]{cc}\frac{(z+3\sigma)(z+\frac{1}{3}\sigma)}{(z-\sigma)(z+\sigma)}&-\frac{(z-\sigma)}{4(z+\sigma)}\\ -\frac{(z-\sigma)}{4(z+\sigma)}&-\frac{3(z-\sigma)}{4(z+\sigma)}\\ \end{array}\right).$ (47)
In this example the central rod is a bolt in the terminology of [14], that is to say it corresponds to an $S^{2}$ in the manifold on which the Killing vector $K=\partial_{\chi}$ vanishes. The central rod of the Riemannian Kerr metric is also a bolt but the central node of the self-dual Taub-NUT solution is a nut, as it is an isolated zero of a Killing vector (in fact of $\partial_{\chi}$).
4.3.2 Hermiticity of Taub-NUT
Both SD and anti-SD Weyl tensors are type D for Taub-NUT, so it has two Hermitian structures, with opposite orientations. With the metric as in (38) consider the basis of 1-forms
$\Theta^{0}=\frac{dR}{\sqrt{U}},\;\Theta^{1}=(R^{2}-N^{2})^{1/2}d\theta,\;\Theta^{2}=(R^{2}-N^{2})^{1/2}\sin\theta d\phi,\;\Theta^{3}=\sqrt{U}(d\chi+2N\cos\theta d\phi),$
then an integrable complex structure is defined by taking the holomorphic 1-forms to be
$\Theta^{0}+i\Theta^{3},\;\;\Theta^{1}+i\Theta^{2}.$
The candidate Kähler form is
$\Omega=\Theta^{0}\wedge\Theta^{3}+\Theta^{1}\wedge\Theta^{2}$
and is not closed but
$\Omega_{K}=(R-N)^{-2}\Omega$
is, so that $(R-N)^{-2}g_{R}$ with $g_{R}$ as in (38) is Kähler.
For the other orientation, switch the sign on $\chi$ and $N$ then the same choices give the other complex structure and now $(R+N)^{-2}g_{R}$ is Kähler.
4.3.3 Conicality
See [14] for a discussion of conicality in self-dual Taub-NUT and the Eguchi-Hanson metric (coming up in Section 4.5); the absence of conical singularities in Taub-bolt is shown in [27].
4.4 Kerr-Taub-bolt
The Kerr-Taub-bolt metric of [15] is in this family, in the sense of having two nodes. Its Weyl tensor is type D, so it’s Hermitian, and it’s easily seen to be AF. Now it isn’t in the list of regular AF Hermitian metrics in [3] so it must not be regular – as also pointed out in [8] it has conical singularities.
Here it is anyway
$g=X\left(\frac{dR^{2}}{\Delta}+d\theta^{2}\right)+\frac{\sin^{2}\theta}{X}(adt+P_{R}d\phi)^{2}+\frac{\Delta}{X}(dt+P_{\theta}d\phi)^{2},$
with
$X=R^{2}-(a\cos\theta+N)^{2}$
$P_{R}=r^{2}-a^{2}-\frac{N^{2}}{N^{2}-a^{2}}$
$P_{\theta}=-a\sin^{2}\theta+2n\cos\theta-\frac{an^{2}}{n^{2}-a^{2}}$
$\Delta=R^{2}-2mR+N^{2}-a^{2}.$
Following the algorithm, calculate
$r=\Delta^{1/2}\sin\theta,\;z=(R-m)\cos\theta$
and then for the Killing vector $K=\partial_{t}$ the twist potential is
$\psi=\frac{2}{X}(NR-m(a\cos\theta+N)).$
Restrict to the upper axis $\theta=0$ and compute from (17) [^8] Previously given in [16].:
$P_{+}=\left(\begin{array}[]{cc}\frac{z^{2}+2mz+m^{2}-(a+N)^{2}}{z^{2}-\sigma^{2}}&\frac{2(Nz-am)}{z^{2}-\sigma^{2}}\\ \frac{2(Nz-am)}{z^{2}-\sigma^{2}}&\frac{-z^{2}+2mz-m^{2}+(a-N)^{2}}{z^{2}-\sigma^{2}}\\ \end{array}\right),$ (48)
where $\sigma^{2}=m^{2}+a^{2}-N^{2}$ and the nodes are at ±$\pm\sigma$. Looking ahead to the inverse problem in Section 5, we note that this matrix actually follows from the determinant condition applied to the two-node ansatz with asymptotic quantities $(m,a,N)$ and free parameters $(A,B,\sigma)$:
$P_{+}=\left(\begin{array}[]{cc}\frac{z^{2}+2mz+A}{z^{2}-\sigma^{2}}&\frac{2(Nz-am)}{z^{2}-\sigma^{2}}\\ \frac{2(Nz-am)}{z^{2}-\sigma^{2}}&\frac{-z^{2}+2mz-m^{2}+B}{z^{2}-\sigma^{2}}\\ \end{array}\right),$ (49)
i.e. this ansatz plus the determinant condition fixes $P_{+}$, including the location of the nodes, from the asymptotics and asymptotic quantities.
4.5 The Gibbons-Hawking metrics
These are characterised as being hyper-Kähler with a triholomorphic Killing vector, or alternatively as Ricci-flat with a self-dual Weyl tensor and a Killing vector whose derivative is self-dual [13, 33]. The metric can be written
$g=V(dx^{2}+dy^{2}+dz^{2})+V^{-1}(dt+{\bf{\omega}}\cdot{\bf{dx}})^{2},$ (50)
in terms of a function $V$ and a 3-vector ${\bf{\omega}}$. The three Kähler forms can be taken to be
$\Omega_{1}=(dt+{\bf{\omega}}\cdot{\bf{dx}})\wedge dx-Vdy\wedge dz$
and its two cyclic permutations, and then closed-ness of these implies the field equation
$\partial_{1}V=\partial_{2}\omega_{3}-\partial_{3}\omega_{2}\mbox{ and cyclic permutations,}$,
or in a 3-vector notation
$\nabla V=\nabla\times{\bf{\omega}}.$
It follows that $V$ is harmonic w.r.t. the flat 3-metric.
It’s worth noting that for some authors (e.g. [6]) a gravitational instanton is necessarily hyper-Kähler.
We are interested in toric symmetry so we need a second Killing vector which commutes with $K=\partial_{t}$. We don’t a priori know what it has to be, so write it in its most general form as
$L=A\partial_{x}+B\partial_{y}+C\partial_{z}+D\partial_{t},$
[…]
$g=V(dr^{2}+r^{2}d\phi^{2}+dz^{2})+V^{-1}(dt+Wd\phi)^{2},$ (51)
with $V(r,z)$ and $W(r,z)$ related by
$rV_{z}=W_{r},\;\;rV_{r}=-W_{z}.$
For $J$ we find
$J=\left(\begin{array}[]{cc}V^{-1}&V^{-1}W\\ V^{-1}W&V^{-1}W^{2}+r^{2}V\\ \end{array}\right),$
[…]
Following the method of Section 2, we can find the twist potential for $K$ as
$\psi=\frac{1}{V}+\mbox{ constant}.$
While it may sometimes be convenient to choose the constant to set $\psi=0$ at infinity, if we take the constant to be zero then from (17) we obtain the simple expression
$P=J^{\prime}(0,z)=\left(\begin{array}[]{cc}V(0,z)&-1\\ -1&0\\ \end{array}\right),$ (52)
and this should hold on all rods (compare (45)).
The two classes of interesting (smooth and ALF or ALE) examples are the multi-Taub-NUT and the multi-Eguchi-Hanson metrics. Both have a $V$ corresponding to a series of point masses all of mass $m$ at a sequence of nodes $z=a_{i},i=1,\ldots,n$ along the $z$-axis, but with $V=1$ at large distances for multi-Taub-NUT and $V=0$ at large distances for multi-Eguchi-Hanson:
$V_{mTN}$ $=$ $1+\sum_{i=1}^{n}\frac{m}{(r^{2}+(z-a_{i})^{2})^{1/2}}$ (53)
$V_{mEH}$ $=$ $\sum_{i=1}^{n}\frac{m}{(r^{2}+(z-a_{i})^{2})^{1/2}}.$ (54)
The top-left entry in $P$ for multi-Taub-NUT is now
$V(0,z)=1+\sum_{i=1}^{n}\frac{m}{(z-a_{i})},$
and the same without the one for multi-Eguchi-Hanson, and as we proceed down the axis from large positive $z$, as we pass the node $a_{i}$ say we simply switch the term $m(z-a_{i})^{-1}$ to $m(a_{i}-z)^{-1}$ in this sum.
Comment: Note that, at a given integer $n$,mTN and mEH have the same arrangement of rods, but different asymptotics and consequently different $P$.
4.5.1 Conicality
All the angular singularities can be removed (see [13]) so these metrics provide regular examples with as many rods as desired, and examples of both ALE (mEH) and ALF (mTN).
4.5.2 Flat space another way
It’s well-known that the Eguchi-Hanson metric in Gibbons-Hawking form but with $n=1$ is actually flat. In this case we can take
$P_{+}=\left(\begin{array}[]{cc}1/z&-1\\ -1&0\\ \end{array}\right),$
which is different from (28): the rod structure here, as in Section 3.1, has a single node, but here the Killing vectors vanishing on the upper and lower rods are the same, while in Section 3.1 they were different. Now consider the transformation
$P_{+}\rightarrow CP_{+}C^{T}\mbox{ with }C=\left(\begin{array}[]{cc}\alpha^{-1}&0\\ 0&\alpha\\ \end{array}\right)\left(\begin{array}[]{cc}1&0\\ \beta z/\alpha&1\\ \end{array}\right)\left(\begin{array}[]{cc}1&0\\ \gamma/\alpha&1\\ \end{array}\right)=\left(\begin{array}[]{cc}\alpha^{-1}&0\\ \beta z+\gamma&\alpha\\ \end{array}\right).$ (55)
Note all three factors of $C$ have determinant one, so det $P_{+}$ is unchanged. The first factor rescales the diagonal entries; the second effects a change in the choice of Killing vector on the upper rod [^9] For the justification of this claim, see the discussion around eqn(9) in [12].; and the third corresponds to adding a constant to the twist potential $\psi$. The effect on $P_{+}$ is then
$P_{+}\rightarrow\left(\begin{array}[]{cc}1/\alpha^{2}z&((\alpha+\beta)z+\gamma)/(\alpha z)\\ ((\alpha+\beta)z+\gamma)/(\alpha z)&((2\alpha\beta+\beta^{2})z^{2}+2\gamma(\alpha+\beta)z+\gamma^{2})/z\\ \end{array}\right).$
If we now take $\gamma=0,\alpha=-\beta=\sqrt{2}$ then $P_{+}$ becomes the $P_{+}$ of (28). What is new here is the middle factor in (55), and this is taken from [12]: the result there is that this is, quite generally, the way to modify $P$ under a change of choice of Killing vector. We’ll make use of it in the discussion of Plebański-Demiański metrics below.
4.5.3 Another particular Gibbons-Hawking metric
This metric was discussed in [16] but originated in [1]. The metric is
$g=f(d\rho^{2}+d\theta^{2})+f^{-1}\sin^{2}\theta(ad\chi+\sigma\cosh\rho d\phi)^{2}+f^{-1}\sinh^{2}\rho(bd\chi+\sigma\cos\theta d\phi)^{2},$ (56)
with
$f=b\cosh\rho-a\cos\theta,$
and I’ve introduced a parameter $\sigma$ not present in [1] (or perhaps set equal to one there, though it is useful to have it present for our purposes). The fact that the metric is in this section implies that the metric is of Gibbons-Hawking type, which wasn’t mentioned in [1] and may not have been known. It was pointed out to me by Maciej Dunajski.
One calculates $J$ and then its determinant to find the Weyl-Papapetrou coordinates as
$r=\sigma\sin\theta\sinh\rho,\;\;z=\sigma\cos\theta\cosh\rho,$
and these are confocal-elliptic-hyperbolic coordinates: the curves of constant $\rho$ are ellipses, the curves of constant $\theta$ are hyperbolas, all with foci at $r=0,z=\pm\sigma$. Also
$z+ir=\sigma\cosh(\rho+i\theta),$
so that
$dz^{2}+dr^{2}=\sigma^{2}(\cosh^{2}\rho-\cos^{2}\theta)(d\rho^{2}+d\theta^{2}).$
Comparing (56) with (50) we conclude that
$V=\frac{f}{\sigma^{2}(\cosh^{2}\rho-\cos^{2}\theta)}=\frac{(\alpha+\beta)\cosh\rho-(\beta-\alpha)\cos\theta}{\sigma^{2}(\cosh^{2}\rho-\cos^{2}\theta)}$
where we’ve put $a=\beta-\alpha,b=\alpha+\beta$, so that
$V=\frac{\alpha}{\sigma(\cosh\rho-\cos\theta)}+\frac{\beta}{\sigma(\cosh\rho+\cos\theta)}.$ (57)
Should we need it, note that the other ingredient in (50) is now
$\omega=Ad\phi\mbox{ with }A=\frac{\alpha(\cosh\rho\cos\theta-1)}{(\cosh\rho-\cos\theta)}+\frac{\beta(\cosh\rho\cos\theta+1)}{(\cosh\rho+\cos\theta)}.$
Putting the expression for $V$ back in terms of $r,z$ we obtain
$V=\frac{\alpha}{(r^{2}+(z-\sigma)^{2})^{1/2}}+\frac{\beta}{(r^{2}+(z+\sigma)^{2})^{1/2}}$
which is the Newtonian potential for point masses $\alpha,\beta$ on the axis at $z=\sigma$ or $-\sigma$ respectively. This is like the Eguchi-Hanson metric with two nodes but with unequal masses. The metric will not be smooth, as already observed in [1], but the corresponding $P_{+}$ will be interesting below. From (52) this is
$P_{+}=\left(\begin{array}[]{cc}\frac{bz-a\sigma}{z^{2}-\sigma^{2}}&-1\\ -1&0\\ \end{array}\right),$ (58)
which we return to in Section 5.
4.6 The Weyl solutions and double Schwarzschild
If we put $\omega=0$ in (5,7,13) then also $\psi=0$ and these are the (Lorentzian) Weyl static solutions. Note that for these $f=e^{2U}$ with $U$ harmonic. As is well-known, the Schwarzschild solution lies in this family with $U$ being the potential for a notional massive rod (which is also a rod in the sense of rod structure) with mass $1/2$ per unit length. It is possible to superimpose any number of distinct but collinear such massive rods to produce a multi-Schwarzschild metric, though one cannot eliminate all the conical singularities, [22].
In the simplest case, for the double Schwarzschild, suppose there are two massive rods lying along the $z$-axis, one of length $b_{1}$ with ends at $a_{0}$ and $a_{1}=a_{0}+b_{1}$ and the other of length $b_{2}$ with ends at $a_{2}$ and $a_{3}=a_{2}+b_{2}$, and we’ll suppose the $a_{i}$ are in order so that $a_{0}<a_{1}<a_{2}<a_{3}$. In terms of rod structure, there are five rods and four nodes at $z=a_{i}$. Where the massive rods are, the Killing vector $\partial_{t}$ is null and on the other three rods $\partial_{\phi}$ vanishes. We’ll label the rods from the bottom as $P_{-},P_{1},P_{2},P_{3},P_{+}$.
Following [22], introduce
$z_{i}=z-a_{i},\;\rho_{i}=(r^{2}+z_{i}^{2})^{1/2},$
then consider
$U=\frac{1}{2}\log\left(\frac{\rho_{0}+\rho_{1}-b_{1}}{\rho_{0}+\rho_{1}+b_{1}}\right)+\frac{1}{2}\log\left(\frac{\rho_{2}+\rho_{3}-b_{2}}{\rho_{2}+\rho_{3}+b_{2}}\right).$ (59)
This is harmonic, and is the $U$ for double Schwarzschild from [22]. These authors also give an explicit expression for $k$, which we don’t need.
At once
$f=e^{2U}=\frac{(\rho_{0}+\rho_{1}-b_{1})(\rho_{2}+\rho_{3}-b_{2})}{(\rho_{0}+\rho_{1}+b_{1})(\rho_{2}+\rho_{3}+b_{2})},$
and we can read off the patching matrices from (13): on the axis with $z>a_{3}$ we have
$\rho_{0}=z-a_{0},\;\rho_{1}=z-a_{1},\;\rho_{2}=z-a_{2},\;\rho_{3}=z-a_{3},$
so that
$f=\frac{(z-a_{1})(z-a_{3})}{(z-a_{0})(z-a_{2})},$ (60)
and
$P_{+}=\left(\begin{array}[]{cc}\frac{(z-a_{0})(z-a_{2})}{(z-a_{1})(z-a_{3})}&0\\ 0&\frac{(z-a_{1})(z-a_{3})}{(z-a_{0})(z-a_{2})}\\ \end{array}\right).$ (61)
In the same way, on the middle rod where $a_{1}<z<a_{2}$ we arrive at
$P_{2}=\left(\begin{array}[]{cc}\frac{(z-a_{0})(z-a_{3})}{(z-a_{1})(z-a_{2})}&0\\ 0&\frac{(z-a_{1})(z-a_{2})}{(z-a_{0})(z-a_{3})}\\ \end{array}\right),$
and on the bottom rod, where $z<a_{0}$
$P_{-}=\left(\begin{array}[]{cc}\frac{(z-a_{1})(z-a_{3})}{(z-a_{0})(z-a_{2})}&0\\ 0&\frac{(z-a_{0})(z-a_{2})}{(z-a_{1})(z-a_{3})}\\ \end{array}\right).$
To obtain $P_{1},P_{3}$ we follow the rule in section 3.4 for going round a node. Taking $P_{+}$ round the node at $a_{3}$ we obtain
$P_{3}=\left(\begin{array}[]{cc}\frac{4(z-a_{0})(z-a_{2})(z-a_{3})}{(z-a_{1})}&0\\ 0&\frac{(z-a_{1})}{4(z-a_{0})(z-a_{2})(z-a_{3})}\\ \end{array}\right),$
and taking $P_{2}$ round the node at $a_{1}$ we obtain
$P_{1}=\left(\begin{array}[]{cc}\frac{4(z-a_{0})(z-a_{1})(z-a_{3})}{(z-a_{2})}&0\\ 0&\frac{(z-a_{2})}{4(z-a_{0})(z-a_{1})(z-a_{3})}\\ \end{array}\right).$
Having obtained the complete set of patching matrices, one can check that passing round the nodes at $a_{0}$ and $a_{2}$ is consistent.
It’s clear that one can obtain an arbitrary even number of nodes this way following the pattern of (60,61) but, as we see next, there will always be conical singularities. This example is interesting as it is not Hermitian, equivalently neither of its Weyl spinors is type D.
4.6.1 Conicality
It’s well-known that, while the horizons are regular (in the Lorentzian case), the conical singularities at the vanishing of $\partial_{\phi}$ cannot all be removed. In the double Schwarzschild case, the periodicity of $\phi$ can be chosen to make the two outer rods nonsingular, when the central rod remains singular and is interpreted as a strut keeping the two black holes apart, or alternatively the periodicity of $\phi$ can be chosen to remove the conical singularity on the central rod when the two outer rods remain singular and are interpreted as ropes holding the two black holes apart.
This discussion was for the Lorentzian metric, but it holds for Riemannian with the usual sign changes.
4.7 C-metric
The C-metric (see e.g. [29]) was considered from the current point of view in [11] where it was seen to have three nodes. The metric is vacuum type D and so admits a valence-two Killing spinor: when made Riemannian, it will be Hermitian and conformal to Kähler in two ways. Following [11] we can take the Lorentzian form of the metric to be
$g_{L}=\frac{1}{A^{2}(x+y)^{2}}\left(F(y)dt^{2}-G(x)d\phi^{2}-\frac{dx^{2}}{G(x)}-\frac{dy^{2}}{F(y)}\right),$
with $G(x)=1-x^{2}-\alpha x^{3},F(y)=-1+y^{2}-\alpha y^{3}$ (so $F(y)=-G(-y)$), and real constants $\alpha,A$ where $m=A/(2\alpha)$ is the mass in the Lorentzian version. We’ll assume that $m$ and $\alpha$ are both nonzero.
We obtain the Riemannian version by replacing $t$ by $it$ and switching the overall sign:
$g_{R}=\frac{1}{A^{2}(x+y)^{2}}\left(F(y)dt^{2}+G(x)d\phi^{2}+\frac{dx^{2}}{G(x)}+\frac{dy^{2}}{F(y)}\right).$ (62)
There are two distinct Hermitian structures defined by taking either $(idt+dy/F,id\phi+dx/G)$ or $(idt+dy/F,id\phi-dx/G)$ as holomorphic one-forms. We note that the 2-forms
$\omega_{\pm}=dt\wedge dy\pm d\phi\wedge dx$
are Kähler w.r.t. one or other of the Hermitian structures for the rescaled metric $(x+y)^{2}g_{R}$ so that also the coordinates $x$ and $y$ are Hamiltonians for the Killing vectors $\partial_{\phi},\partial_{t}$ respectively.
The metric functions $F$ and $G$ both have 3 real roots if $\alpha^{2}<4/27$, which we’ll assume. Label the roots of $G$ as $\beta_{1}<\beta_{2}<0<\beta$, then the roots of $F$ are $-\beta<0<-\beta_{2}<-\beta_{1}$ (also we can write $G(x)=-\alpha(x-\beta)(x-\beta_{1})(x-\beta_{2})$ or $F=-\alpha(y+\beta)(y+\beta_{1})(y+\beta_{2})$ whenever desired). For a Riemannian metric we need $F(y)$ and $G(x)$ non-negative and there are four regions in the $(x,y)$-plane that would satisfy this. However, for one of them the region doesn’t meet the line $x+y=0$, which represents infinity (as is evident from (62)), and in two more of them $x$ or $y$ can go to infinity (and the metric is singular), so we’ll choose the fourth, which is the region $R$ defined by $\beta_{2}\leq x\leq\beta$ and $-\beta_{2}\leq y\leq-\beta_{1}$. Note that $F=0$ on two sides of this region and $G=0$ on the other two: the coordinate $r$ will vanish all the way round the rectangle.
The Weyl-Papapetrou coordinates are [11]
$r=\frac{(F(y)G(x))^{1/2}}{A^{2}(x+y)^{2}},\;\;z=\frac{(\alpha xy^{2}-\alpha x^{2}y-2xy-2)}{2A^{2}(x+y)^{2}}.$
If we confine $(x,y)$ to the rectangle $\beta_{2}<x<\beta,-\beta_{2}<y<-\beta_{1}$, then the corner $(\beta_{2},-\beta_{2})$ is at infinity and the other three corners are nodes, so the four sides are the rods, and we see at once which Killing vector vanishes on each rod: $\partial_{t}$ when $F=0$ and $\partial_{\phi}$ when $G=0$.
Following [11] we can calculate the $z$-coordinates of the three nodes to find
$z(\beta,-\beta_{2})=\mu:=\frac{\alpha\beta_{1}}{2A^{2}},\;z(\beta,-\beta_{1})=\nu:=\frac{\alpha\beta_{2}}{2A^{2}},\;z(\beta_{2},-\beta_{1})=\xi:=\frac{\alpha\beta}{2A^{2}}$
with $\mu<\nu<0<\xi$. Since the Killing vectors are both twist-free, all the patching matrices will be diagonal (see (17)) and we just need to restrict $g_{tt}$ or $g_{\phi\phi}$ to the axis. On the top-most rod $z>\xi$ the Killing vector $\partial_{\phi}$ vanishes so we use $\partial_{t}$ to construct the patching matrix $P_{+}$ according to (17), and we simply need
$f=g_{tt}=\frac{F(y)}{A^{2}(\beta_{2}+y)^{2}}=\frac{-\alpha(y+\beta)(y+\beta_{1})}{A^{2}(y+\beta_{2})}$
where we have substituted the $x$-value for this rod, namely $\beta_{2}$, and used the factorisation of $F(y)$. We want this in terms of $z$ where
$z=\frac{(\alpha\beta_{2}y^{2}-\alpha\beta_{2}^{2}y-2\beta_{2}y-2)}{2A^{2}(\beta_{2}+y)^{2}}=\frac{(\alpha\beta_{2}^{2}y-2)}{2A^{2}\beta_{2}(y+\beta_{2})}.$
Eliminating $y$ from $f$ in favour of $z$ we find the simple expression
$f=\frac{2(z-\mu)(z-\xi)}{(z-\nu)}.$ (63)
For the other three rods we employ the switching procedure of subsection 3.4. The function $f$ above is replaced in turn by
$\frac{(z-\mu)}{2(z-\xi)(z-\nu)},\;\frac{2(z-\mu)(z-\nu)}{(z-\xi)},\;\mbox{ and}\frac{(z-\nu)}{2(z-\mu)(z-\xi)},$
as one passes the nodes at $\xi,\nu,\mu$ respectively. By the end, on the bottom rod, we’ve arrived at the inverse $P$ of the top rod. This can be compared with $\mathbb{E}^{4}$ in subsection 4.1 and contrasted with the Kerr case in subsection 4.2 and the double Schwarzschild in subsection 4.4 where the patching matrix $P_{+}$ for the top rod is the same as the patching matrix $P_{-}$ for the bottom rod. This is because the asymptotics are different: Kerr and double Schwarzschild are AF while this C-metric is AE. To see this for the Riemannian C-metric, we look near to the point $(x,y)=(\beta_{2},-\beta_{2})$: set
$x=\beta_{2}+\frac{c}{R^{2}}\cos^{2}(\theta/2),\;\;y=-\beta_{2}+\frac{c}{R^{2}}\sin^{2}(\theta/2),$
where $c$ is a constant to be chosen, and we’re thinking of $R$ as large. Then
$G(x)=k\frac{c}{R^{2}}\cos^{2}(\theta/2),\;F(y)=k\frac{c}{R^{2}}\sin^{2}(\theta/2)$
where $k=G^{\prime}(\beta_{2})=F^{\prime}(-\beta_{2})$ which can be seen from the graph of $G$ to be positive, and we neglect higher order in $1/R$. Substitute for $F$ and $G$ into (62), along with $t=\alpha-\beta,\phi=\alpha+\beta$ to find
$g_{R}=\frac{4k}{A^{2}c}\left(dR^{2}+\frac{1}{4}R^{2}(d\theta^{2}+\sin^{2}\theta d\beta^{2}+(d\alpha+\cos\theta d\beta)^{2}))\right)$
which is the standard flat metric on $\mathbb{R}^{4}$ if we set $c=4k/A^{2}$.
Being static and axisymmetric, the C-metric is in Weyl’s class and arises from a harmonic potential $U$ similar in form to that in (59) but with an odd number of summands. Following the pattern of the double-Schwarzschild, we may write down a $P$ with any odd number of nodes, for example for nodes at $z=a_{1},\ldots,a_{5}$ take
$f:=\frac{2(z-a_{1})(z-a_{3})(z-a_{5})}{(z-a_{2})(z-a_{4})},$
in place of (60) and (63). Again it will not be possible to remove all the conical singularities, but the metric will this time be ALE and still not Hermitian.
There is a generalisation of the C-metric known as the spinning or twisting C-metric (see e.g. [28]). The metric is not diagonal and so neither is $P$ but this metric is a special case of the Plebański-Demiański metrics which in turn are a special case of the Chen-Teo metric so we’ll leave it until we look at those.
4.7.1 Conicality
The conical singularities on the axis cannot all be resolved (a familiar fact in the Lorentzian case where they are sometimes interpreted as struts holding two black holes apart). To see this, consider the rod at $x=\beta$ and calculate the periodicity in $\phi$ needed to make this regular: nearby $x=\beta-\rho^{2}$ so $G(x)=-G^{\prime}(\beta)\rho^{2}+h.o.$ and the condition is
$\Delta\phi=4\pi/(-G^{\prime}(\beta)),$
while for the rod at $x=\beta_{2}$ one needs
$\Delta\phi=4\pi/(G^{\prime}(\beta_{2})),$
and these cannot be the same:
$G^{\prime}(\beta)+G^{\prime}(\beta_{2})=-\alpha(\beta-\beta_{2})^{2}\neq 0.$
Similarly the rods at $y=-\beta_{1},-\beta_{2}$ require
$\Delta t=4\pi/(-G^{\prime}(\beta_{1}))\mbox{ or }4\pi/(G^{\prime}(\beta_{2})),$
which are incompatible:
$G^{\prime}(\beta_{1})+G^{\prime}(\beta_{2})=-\alpha(\beta_{1}-\beta_{2})^{2}\neq 0.$
4.8 Plebański-Demiański metrics
These metrics, [28], provide a useful step on the way to the Chen-Teo metric. In fact in [8] Chen and Teo describe their new metric as a bridge between Plebański-Demiański and a 3-node Gibbons-Hawking metric. They can be Lorentzian or Riemannian and are still type-D (or Hermitian) but in the Riemannian case now ALE not ALF. A cosmological constant can be included, but we’ll omit that.
Take the Riemannian version of the metric as given in [8]:
$g=A^{2}dp^{2}+B^{2}dq^{2}+(\alpha d\tau+\beta d\phi)^{2}+(\lambda d\tau+\mu d\phi)^{2},$ (64)
where
$A^{2}=\frac{(1-p^{2}q^{2})}{(p-q)^{2}P(p)},\;B^{2}=-\frac{(1-p^{2}q^{2})}{(p-q)^{2}Q(q)}$
$\alpha^{2}=\frac{q^{4}P}{(p-q)^{2}(1-p^{2}q^{2})},\;\beta^{2}=\frac{P}{(p-q)^{2}(1-p^{2}q^{2})},\;\alpha\beta<0$
$\lambda^{2}=-\frac{Q}{(p-q)^{2}(1-p^{2}q^{2})},\;\mu^{2}=-\frac{p^{4}Q}{(p-q)^{2}(1-p^{2}q^{2})},\;\lambda\mu<0$
and
$P(p)=\gamma+2np-\epsilon p^{2}+2mp^{3}+\gamma p^{4},\;Q(q)=\gamma+2nq-\epsilon q^{2}+2mq^{3}+\gamma q^{4}.$ (65)
Although $P,Q$ are the same function, we’ll adopt the convention that $P$ always means $P(p)$ and $Q$ always means $Q(q)$. There will be space-time singularities where $p^{2}q^{2}=1$, and points with $p=q$ must be at infinity. Very roughly, we can think of $m$ and $n$ as mass and nut-parameter, but $\epsilon$ is not charge (since the solution is vacuum).
For the signature to be Riemannian, we need $P$ positive and $Q$ negative (w.l.o.g. we can suppose $\gamma>0$). Thus, similarly to the $C$-metric, we need $p$ to lie between two roots of $P$ in an interval, say $I_{1}$, with $P$ positive, and $q$ to lie between two roots of $Q$ in an interval, say $I_{2}$, with $Q$ negative. We’ll suppose that $P$ has four real roots, ordered $-\infty<p_{1}<p_{2}<0<p_{3}<p_{4}<\infty$ so there are three possibilities for $I_{1}$, namely $(-\infty,p_{1}),(p_{2},p_{3})$ or $(p_{4},\infty)$, and two possibilities for $I_{2}$, namely $(p_{1},p_{2})$ or $(p_{3},p_{4})$, giving six possible rectangles altogether. The rectangles with a semi-infinite side can be ruled out as inevitably they will contain points with $p^{2}q^{2}=1$. Therefore we want $p_{2}\leq p\leq p_{3}$ and either $p_{1}\leq q\leq p_{2}$ or $p_{3}\leq q\leq p_{4}$. Clearly points with $p=q$ are at infinity and one corner of either choice of the rectangle $I_{1}\times I_{2}$ lies on the line $p=q$, and is at infinity. The other corners give three nodes. The constants in $P$ must be chosen so that the singularities which occur at $p^{2}q^{2}=1$ don’t meet the chosen rectangle, which we’ll assume can be arranged for at least one of the two rectangles.
With the conventions of (1) we have
$U=\alpha^{2}+\lambda^{2},\;V=\alpha\beta+\lambda\mu,\;W=\beta^{2}+\mu^{2},$
so that
$r^{2}=UW-V^{2}=(\alpha\mu-\beta\lambda)^{2}$
from which we may take
$r=\frac{(-PQ)^{1/2}}{(p-q)^{2}}.$ (66)
The axis is now at $PQ=0$ and the rods are at zeroes of $P$ or of $Q$ with nodes at common zeroes of both.
After a calculation we find the function conjugate to $r$, up to an additive constant, is
$z=(-\gamma-n(p+q)+\epsilon pq-\gamma p^{2}q^{2}-mpq(p+q))/(p-q)^{2}.$ (67)
On a rod or section of the axis where $P=0$ we may add $P$ to the numerator of $z$ to find the simplification
$z=[(n-\epsilon p+2mp^{2}+\gamma p^{3})+(mp+\gamma p^{2})q]/(p-q),$
so that, since $p$ is constant on this rod, $z$ is a Möbius fiunction of $q$. Furthermore along this rod
$\frac{dz}{dq}=\frac{1}{2(p-q)^{2}}\frac{dP}{dp}.$
There is a similar conclusion on a rod with $Q=0$, when $z$ is a Möbius function of $p$:
$z=[(-n+\epsilon q-2mq^{2}-\gamma q^{3})-(mq+\gamma q^{2})p]/(p-q),$
and
$\frac{dz}{dp}=\frac{1}{2(p-q)^{2}}\frac{dQ}{dq}.$
From (66) we see that there are nodes at the common zeroes of $P$ and $Q$ so these will be at the corners of $I_{1}\times I_{2}$ and there are two possibilties:
  • going clockwise round the rectangle from $(p_{3},p_{3})$, the point at infinity, the nodes are at
    $(p,q)=(p_{2},p_{3}),\;(p_{2},p_{4}),\;(p_{3},p_{4}),$
  • with corresponding $z$-values
    $z_{1}=-\frac{\epsilon}{2}+m(p_{2}+p_{3})+\frac{\gamma}{2}(p_{2}+p_{3})^{2},$
    $z_{2}=-\frac{\epsilon}{2}+m(p_{2}+p_{4})+\frac{\gamma}{2}(p_{2}+p_{4})^{2},$
    $z_{3}=-\frac{\epsilon}{2}+m(p_{3}+p_{4})+\frac{\gamma}{2}(p_{3}+p_{4})^{2},$
  • by inspection of $dz/dq$ and $dz/dp$ along the rods, one sees that clockwise round the rectangle is the sense of increasing $z$ so the top rod is $p=p_{3}$ and the bottom rod is $q=p_{3}$,
or
  • going clockwise round the rectangle from $(p_{2},p_{2})$, the point at infinity, the nodes are at
    $(p,q)=(p_{3},p_{2}),\;(p_{3},p_{1}),\;(p_{2},p_{1}),$
  • with corresponding $z$-values
    $z_{1}=-\frac{\epsilon}{2}+m(p_{3}+p_{2})+\frac{\gamma}{2}(p_{3}+p_{2})^{2},$
    $z_{2}=-\frac{\epsilon}{2}+m(p_{3}+p_{1})+\frac{\gamma}{2}(p_{3}+p_{1})^{2},$
    $z_{3}=-\frac{\epsilon}{2}+m(p_{2}+p_{1})+\frac{\gamma}{2}(p_{2}+p_{1})^{2},$
  • by inspection of $dz/dq$ and $dz/dp$ along the rods, again clockwise round the rectangle is the sense of increasing $z$ so the top rod is $p=p_{2}$, and the bottom rod is $q=p_{2}$.
After some work, we obtain the twist potential for $K=\partial\tau$ by solving
$d\psi=*(K\wedge dK)$
as
$\psi=\frac{F(p,q)}{(p-q)(1-p^{2}q^{2})}$ (68)
with
$F=-2\gamma q-2nq^{2}+2mpq^{3}+2\gamma p^{2}q^{3}.$ (69)
Now we have enough to write down $P_{+}$ but the details are messy and will appear elsewhere. It’s fairly straightforward to get the algebraic form as
$P_{+}=\left(\begin{array}[]{cc}\frac{q_{1}}{C_{1}}&\frac{C_{2}}{C_{1}}\\ \frac{C_{2}}{C_{1}}&\frac{Q_{1}}{C_{1}}\\ \end{array}\right),$
where $q_{1}$ is quadratic in $z$, $C_{1},C_{2}$ are both cubic and $Q_{1}$ is quartic. Recall these metrics are AE/ALE but still this is not in the form that one expects from (26). However one has the transformation in (55) available and this can be used to remove the two highest terms in $C_{2}$, leaving it linear, say
$P_{+}=\left(\begin{array}[]{cc}\frac{q_{1}}{C_{1}}&\frac{\ell}{C_{1}}\\ \frac{\ell}{C_{1}}&\frac{\widetilde{Q}_{1}}{C_{1}}\\ \end{array}\right),$ (70)
with linear $\ell(z)$, and $\widetilde{Q}_{1}$ a different quartic from the previous $Q_{1}$. We’ll return to this form for $P_{+}$ in Section 5.3.
4.8.1 Conicality
There’s no hope of removing all the conical singularities. The four rods have four different constant combinations of $\partial_{\tau},\partial_{\phi}$ vanishing, in order from the bottom
$p_{3}^{2}\partial_{\tau}-\partial_{\phi},\;\partial_{\tau}-p_{4}^{2}\partial_{\phi},\;p_{2}^{2}\partial_{\tau}-\partial_{\phi},\;\partial_{\tau}-p_{3}^{2}\partial_{\phi}.$
4.9 The Chen-Teo metric
Following [8], the metric is
$g=\frac{kH}{(x-y)^{3}}\left(\frac{dx^{2}}{X}-\frac{dy^{2}}{Y}-\frac{XY}{kF}d\phi^{2}\right)+\frac{(Fd\tau+Gd\phi)^{2}}{(x-y)HF},$ (71)
where
$X=a_{0}+a_{1}x+a_{2}x^{2}+a_{3}x^{3}+a_{4}x^{4}$ (72)
$Y=a_{0}+a_{1}y+a_{2}y^{2}+a_{3}y^{3}+a_{4}y^{4}$ (73)
$F=x^{2}Y-y^{2}X$ (74)
$H=(\nu x+y)[(\nu x-y)(a_{1}-a_{3}xy)-2(1-\nu)(a_{0}-a_{4}x^{2}y^{2})]$ (75)
$G=(\nu^{2}a_{0}+2\nu a_{3}y^{3}+2\nu a_{4}y^{4}-a_{4}y^{4})X+(a_{0}-2\nu a_{0}-2\nu a_{1}x-\nu^{2}a_{4}x^{4})Y.$ (76)
At this point there are seven parameters: $k$ which we mentioned in Section 2.1, $\nu$, and the five coefficients $a_{0},\ldots,a_{4}$ of the polynomials $X$ and $Y$. We assume that $a_{4}>0$ and that $X$ (and therefore also $Y$) has four real roots, $x_{i},i=1,\ldots,4$ with
$-\infty<x_{1}<x_{2}<0<x_{3}<x_{4}<\infty$
and that $x$ is confined to an interval with $X$ positive, while $y$ is confined to an interval with $Y$ negative. As with the Plebański-Demiański metric in the previous section, this leads to two possible rectangles in the $(x,y)$-plane, each with one vertex on the line $x=y$, which will be the point at infinity. In [8] we learn that we must take the rectangle $\{x_{2}\leq x\leq x_{3},\;x_{1}\leq y\leq x_{2}\}$ to obtain a solution without curvature singularities (though conical singularities have yet to be eliminated).
For the Weyl-Papapetrou coordinates, from [8]
$\rho=\frac{(-XY)^{1/2}}{(x-y)^{2}},\;z=\frac{2(a_{0}+a_{2}xy+a_{4}x^{2}y^{2})+(x+y)(a_{1}+a_{3}xy)}{2(x-y)^{2}},$ (77)
so that the axis is at $XY=0$, which is the boundary of whichever coordinate rectangle is chosen, and there are three nodes corresponding to the three corners of this rectangle not at infinity.
On $y=x_{2}$ claim
$z=((a_{1}+2a_{2}x_{2}+2a_{3}x_{2}^{2}+2a_{4}x_{2}^{3})+x(a_{3}x_{2}+2a_{4}x_{2}^{2}))(2(x-x_{2}))^{-1},$
which is Möbius in $x$ and positive, tending to infinity as $x$ tends from above to $x_{2}$, so this is the top rod.
On $x=x_{2}$ claim
$z=((a_{1}+2a_{2}x_{2}+2a_{3}x_{2}^{2}+2a_{4}x_{2}^{3})+y(a_{3}x_{2}+2a_{4}x_{2}^{2}))(2(y-x_{2}))^{-1},$
which is Möbius in $y$ and negative, tending to minus infinity as $y$ tends from below to $x_{2}$, so this is the bottom rod.
The topmost node is where the top rod arrives at $x_{3}$ so this is at
$z=z_{3}=-\frac{1}{2}(a_{2}+a_{3}(x_{2}+x_{3})+a_{4}(x_{2}+x_{3})^{2}).$
The bottom-most node is at $(x,y)=(x_{2},x_{1})$ so that
$z=z_{1}=-\frac{1}{2}(a_{2}+a_{3}(x_{1}+x_{2})+a_{4}(x_{1}+x_{2})^{2}).$
(and it’s easily checked that $z_{3}>z_{1}$).
For the Killing vector $K=\partial_{\tau}$ the twist potential $\psi$ solving
$d\psi=*(K\wedge dK)$
can be taken to be
$\psi=H^{-1}[a_{0}y+(2\nu-1)a_{0}x+(a_{1}\nu-a_{4}y^{3})x^{2}+(-\nu a_{3}y-2\nu a_{4}y^{2}+a_{4}y^{2})x^{3}].$ (78)
We’ve done enough to obtain $J^{\prime}$ and $P_{+}$ but the details are complicated and can be found in [9]. The basic shape, given ALF and three nodes, is
$P_{+}=\Delta^{-1}\left(\begin{array}[]{cc}C_{1}&\ell\\ \ell&C_{2}\\ \end{array}\right),$ (79)
with nodes at $z=a_{1},a_{2},a_{3}$, $\Delta(z)=(z-a_{1})(z-a_{2})(z-a_{3})$, $C_{i}(z)$ cubic and $\ell(z)$ linear. Again we’ll revisit this in Section 5.3.
5 The inverse problem
This is the problem of obtaining $P$ from the data, i.e. from the rod structure and the asymptotics, rather than from the metric. It was solved for four-dimensional AF with 2 nodes in [11, 12] and for some five-dimensional cases in [32].
Let us first note some general properties of $P(z)$: it can’t have a double pole off the axis because this would be a double zero in $f$, i.e. $f$ and $df$ with a common zero which forces the Killing vector $K$ to vanish everywhere; a double pole of $P$ on the axis gives a space-time singularity because the splitting process will fail there; by the splitting process, a simple pole in $P$ off the axis or at a complex $z$ also gives rise to a space-time singularity. We conclude that the only singularities in $P$ are simple poles at the nodes. If these are located at $z=a_{i},i=1,\ldots,n$ then every entry in $P$ is a rational function with denominator
$\Delta:=\Pi_{1}^{n}(z-a_{i}).$
Furthermore the degrees of the numerators are fixed by the asymptotic conditions.
I’ve made no mention of this, but it is known how to recognise (and therefore attempt to remove) conical singularities in this formalism: see p121 in [11].
If we can deduce $P(z)$ from the data then the metric is obtained by splitting $P(z)$. Solving this inverse problem was part of the motivation of [11, 12] as this was seen as a new approach to black hole uniqueness, and indeed it works best with the Kerr solution. The general technique motivates the ansatz for AF/ALF
$P_{+}=\frac{1}{\Delta}\left(\begin{array}[]{cc}P_{1}(z)&Q(z)\\ Q(z)&P_{2}(z)\\ \end{array}\right),$
where $P_{1},P_{2}$ are polynomials of degree $n$ with leading coefficients $+1$ and $-1$ respectively and next coefficients $2M$ for both, and $Q(z)$ is a polynomial of degree $n-1$ with leading coefficient $2N$ and next coefficient $-2L$, while for ALE we take the ansatz
$P_{+}=\frac{1}{\Delta}\left(\begin{array}[]{cc}P(z)&Q(z)\\ Q(z)&S(z)\\ \end{array}\right),$
where $P(z)$ is a polynomial of degree $n-1$ with leading coefficient $1/2$ and next coefficient $-M/2$, $Q(z)$ is a polynomial of degree $n-2$ with leading coefficient $L/2$, and $S(z)$ is a polynomial of degree $n+1$ with leading coefficient $-2$ and next coefficient $-2M$. Then the unit determinant condition will fix more of the free constants.
We’ll order the discussion by the number of nodes. Zero nodes leads to flat space in the AF/ALF case and is inconsistent with AE/ALE so the simplest case to start with is a single node.
5.1 One node
For AF/ALF this case has only the SDTN from (45):
$P_{+}=\frac{1}{z}\left(\begin{array}[]{cc}z+2m&-1\\ -1&0\\ \end{array}\right),$
which we can transform according to (52) with
$C=\left(\begin{array}[]{cc}1&0\\ 1&1\\ \end{array}\right),$
to obtain
$P_{+}=\frac{1}{z}\left(\begin{array}[]{cc}z+2m&2m\\ 2m&-z+2m\\ \end{array}\right),$
and this takes the form of the ansatz but with $N=m$, and is the unique solution.
For AE/ALE we only have $P_{+}$ as in (28), and correspondingly the flat metric on $\mathbb{E}^{4}$.
5.2 Two nodes
We’ll always place the nodes at $z=\pm\sigma$.
Then for AF/ALF the ansatz is
$P_{+}=\frac{1}{z^{2}-\sigma^{2}}\left(\begin{array}[]{cc}z^{2}+2mz+A&2Nz-2L\\ 2Nz-2L&-z^{2}+2mz+B\\ \end{array}\right).$
The determinant condition branches into two cases:
  • $m^{2}=N^{2}$: then the metric takes the Gibbons-Hawking form and the potential is for unequal masses at $z=\pm\sigma$. The solution isn’t regular unless the masses are equal, when it’s the two-centre SDTN.
  • $m^{2}\neq N^{2}$: this is Kerr-Taub-bolt as in (49); in particular it includes Kerr when $N=0$. This is Black-hole Uniqueness: two nodes, Lorentzian and AF implies Kerr.
With AE/ALE and two nodes, we can locate the nodes at $z=\pm\sigma$ as before and the asymptotics (26) force the ansatz:
$P_{+}=\frac{1}{(z-\sigma)(z+\sigma)}\left(\begin{array}[]{cc}\ell(z)&-2L\\ -2L&c(z)\\ \end{array}\right),$ (80)
where $\ell(z)$ is linear and $c(z)$ is cubic with some coefficients fixed by (26):
$\ell(z)=\frac{1}{2}(z-M),\;\;c(z)=-(2z^{3}+2Mz^{2}+2Az+2B),$
where $M,L$ are the asymptotic quantities as in (26). The constants $A,B$ are to be fixed by the determinant condition. This gives
$A=M^{2}-2\sigma^{2},\;\;B=MA,\;\;\sigma^{2}=M^{2}\pm 2L,$
where we have solved a quadratic for $\sigma^{2}$ (which must be positive of course).
It turns out that this $P_{+}$ arises by appying the transformation in (55) to the matric $P_{+}$ in (58): take
$\alpha=(2b)^{1/2},\;\beta=(2/b)^{1/2},\;\gamma=a\sigma(2/b^{3})^{1/2},$
then $P_{+}$ from (58) precisely becomes $P_{+}$ of (80) with
$M=a\sigma/b,\;L=\frac{\sigma^{2}(a^{2}-b^{2})}{2b^{2}}.$
Therefore the general ALF two-node solution is the Bazaĭkin solution (58), which includes the Eguchi-Hanson metric.
5.3 Three nodes
Now for three nodes, suppose they are at $z=\alpha,\beta$ and $\gamma$, where w.l.o.g. $\alpha+\beta+\gamma=0$, and set
$\Delta=(z-\alpha)(z-\beta)(z-\gamma).$
The ansatz for an AF/ALF solution is
$P_{+}(z)=\Delta^{-1}\left(\begin{array}[]{cc}C_{1}(z)&\ell(z)\\ \ell(z)&C_{2}(z)\\ \end{array}\right),$
with
$C_{1}(z)=z^{3}+2Mz^{2}+Az+B,$
$\ell(z)=2L(z-C)$
$C_{2}(2)=-z^{3}+2Mz^{2}+Dz+E,$
with the determinant condition
$C_{1}(z)C_{2}(z)-(q(z))^{2}=-\Delta^{2}.$
This is analysed in [9]. It doesn’t seem to lead precisely to Chen-Teo, and the possibility remains that there are more solutions here. Because of [3], if there are then they won’t be Hermitian.
The ansatz for an ALE solution is
$P_{+}(z)=\Delta^{-1}\left(\begin{array}[]{cc}q(z)&\ell(z)\\ \ell(z)&Q(z)\\ \end{array}\right),$
with
$q(z)=\frac{1}{2}(z^{2}+2Mz+2A),$
$\ell(z)=2L(z-B)$
$Q(2)=-2(z^{4}-2Mz^{3}+Cz^{2}+Dz+E),$
and $A,B,C,D,E$ free. The determinant condition is
$q(z)Q(z)-(\ell(z))^{2}=-\Delta^{2}.$
It can be solved for all quantities in terms of $A$, leaving a quartic for $A$. This case must include the Plebański-Demiański metrics, and the 3-centre Eguchi-Hansen metric with unequal masses.
Appendix 1: Things going wrong
Double poles
As noted above, there can’t be a double pole in $P$ away from the axis because this would be a double zero in $f$, i.e. $f$ and $df$ with a common zero which forces the Killing vector $K$ to vanish everywhere.
One can construct Gibbons-Hawking examples with a double pole on the axis e.g. take $V$ for a simple dipole:
$V=\frac{z}{(r^{2}+z^{2})^{3/2}}$
so that $V(0,z)=z^{-2}$ and $P$ in (52) has a double pole at the origin. However the 4-metric is singular there: anything but a simple pole in $V$ gives a space-time singularity (as follows from the splitting process).
[…]
$P_{+}=\left(\begin{array}[]{cc}\frac{Q(z)}{p^{2}(z^{2}-1)^{2}}&\frac{8qz^{2}}{p^{2}(z^{2}-1)^{2}}\\ \frac{8qz^{2}}{p^{2}(z^{2}-1)^{2}}&\frac{p^{4}(z^{2}-1)^{4}+64q^{2}z^{4}}{p^{2}(z^{2}-1)^{2}Q(z)}\\ \end{array}\right),$
with
$Q(z)=p^{2}(z^{2}-1)^{2}+4zp(z-1)^{2}+8z^{2}(p+1).$
This has AF asymptotics with $m=2/p,L=-4q/p^{2}$, but it has double poles at $z=1$ and extra poles where $Q$ vanishes (which are poles in $f=g_{tt}$; evidently these are space-time singularities if at real $z$, otherwise they are complex poles in $P$, and give rise to space-time singularities from the splitting process).
Appendix 2: Splitting
This is an important part of the Ward construction [35] and it has a particular form in this special case with two symmetries – see [36], [12]. These two references differ in their details so I’m going to follow the latter and give just an abbreviated account of it. I’ll start with a motivating example, which is classical [^10] In the sense of old enough not to need a reference.: suppose $F(z)$ is holomorphic in some part of the complex $z$-plane including at least part of the real axis. Substitute $z-\frac{1}{2}r(\zeta-\zeta^{-1})$ for $z$ and expand $F$ as a Laurent series in $\zeta$:
$F(z-\frac{1}{2}r(\zeta-\zeta^{-1}))=\Sigma_{-\infty}^{\infty}a_{n}(r,z)\zeta^{n},$
then differentiating the series term by term we find that
$\Sigma_{-\infty}^{\infty}[ra_{n,rr}+a_{n,r}+ra_{n,zz}-\frac{n^{2}}{r}a_{n}]\zeta^{n}=0.$
In particular therefore $a_{0}(r,z)$ is harmonic and this term can be extracted by a contour integral:
$F(r,z):=a_{0}(r,z)=\frac{1}{2\pi i}\oint F(z-\frac{1}{2}r(\zeta-\zeta^{-1}))\frac{d\zeta}{\zeta}.$
Furthermore, evaluating at $r=0$ we find $F(0,z)=F(z)$ so this is the harmonic function with value $F(z)$ on the axis.
We connect to the patching matrix following [12]: with patching matrix $P(z)$ substitute for $z$ as before and split by factorising:
$P(z-\frac{1}{2}r(\zeta-\zeta^{-1}))=H(r,z;\zeta)\widehat{H}^{-1}(r,z;\zeta),$
where $\widehat{H}(r,z;\zeta)$ is holomorphic in $\zeta$ for $|\zeta|\geq 1$, including $\infty$, and $H(r,z;\zeta)$ is holomorphic in $\zeta$ for $|\zeta|\leq 1$ (this is always possible). Now set
$J^{\prime}(r,z)=H(r,z;0)\widehat{H}^{-1}(r,z;\infty),$ (81)
then this is the desired Ernst potential from $P(z)$ and satisfies the Yang equation (11).
For a simple example, consider (17) with a static metric, so that $\psi=0$ and by (18)
$P(z)=\left(\begin{array}[]{cc}(f(0,z))^{-1}&0\\ 0&f(0,z)\\ \end{array}\right),$
and replace $z$ by $z-\frac{1}{2}r(\zeta-\zeta^{-1})$. Evidently $H,\widehat{H}$ can be taken to be diagonal:
$\widehat{H}=\left(\begin{array}[]{cc}e^{\hat{h}}&0\\ 0&e^{-\hat{h}}\\ \end{array}\right),\;H=\left(\begin{array}[]{cc}e^{-h}&0\\ 0&e^{h}\\ \end{array}\right)$
and so we need
$\hat{h}(r,z;\zeta)+h(r,z;\zeta)=\log f(z-\frac{1}{2}r(\zeta-\zeta^{-1})),$
which we solve by expanding the r.h.s. as a Laurent series in $\zeta$, as before. Call the $O(0)$ term $a_{0}(r,z)$ then the prescription (81) gives
$J^{\prime}=\left(\begin{array}[]{cc}e^{-a_{0}(r,z)}&0\\ 0&e^{a_{0}(r,z)}\\ \end{array}\right).$
We recover $J(r,z)$ from (15) and the metric by solving (9,10) with $\omega=0$. These are the Weyl solutions – see Section 4.6 or [29].
We can use the same trick to split the patching matrix for the Gibbons-Hawking solutions (52): put $z-\frac{1}{2}r(\zeta-\zeta^{-1})$ for $z$ and consider
$P(z-\frac{1}{2}r(\zeta-\zeta^{-1}))=\left(\begin{array}[]{cc}F(z-\frac{1}{2}r(\zeta-\zeta^{-1}))&-1\\ -1&0\\ \end{array}\right)=H(r,z;\zeta)\widehat{H}^{-1}(r,z;\zeta),$
where $F(z)=V(0,z)$. A shrewd choice is
$H(r,z;\zeta)=\left(\begin{array}[]{cc}1&h\\ 0&-1\\ \end{array}\right),\;\;\widehat{H}(r,z;\zeta)=\left(\begin{array}[]{cc}0&1\\ -1&\hat{h}\\ \end{array}\right),$
and then we just need
$F(z-\frac{1}{2}r(\zeta-\zeta^{-1}))=\hat{h}+h$
so that, with the Laurent series $F=\Sigma_{-\infty}^{\infty}a_{n}(r,z)\zeta^{n}$, we may take
$\hat{h}=\Sigma_{-\infty}^{-1}a_{n}(r,z)\zeta^{n},\;\;h=\Sigma_{0}^{\infty}a_{n}(r,z)\zeta^{n},$
and obtain
$J^{\prime}(r,z)=H(r,z;0)\widehat{H}^{-1}(r,z;\infty)=\left(\begin{array}[]{cc}a_{0}(r,z)&-1\\ -1&0\\ \end{array}\right),$
which is correct since w.l.o.g. $\psi=f=V^{-1}$ in (16), so that $V(r,z)=a_{0}(r,z)$ is harmonic.
A general splitting algorithm is under development by Mason and collaborators.